QUESTION 10 Calculate the value of the error (anly) with one decimal place for (20+1-7%) x (34 +1-8%) Please enter the answer without +- sign.QUESTION 6 Calculate the value of the error (only) with two decimal places for Please enter the answer without +- sign.
Answer #1

(10) Solution:

Here;

dr / r = 0.07 + 0.08= 0.15

=> dr / [ 34 * 20 ] = 0.15

=> dr = 102

Error = 102